# Robin: a Suite of Multi-Scale Vision-Language Models and the CHIRP Evaluation Benchmark

### v2

@@ Abstract
The proliferation of Vision-Language Models (VLMs) in the past several years calls for rigorous and comprehensive evaluation methods and benchmarks. This work analyzes existing VLM evaluation techniques, including automated metrics, AI-based assessments, and human evaluations across diverse tasks. We first introduce Robin - a novel suite of VLMs that we built by combining Large Language Models (LLMs) and Vision Encoders (VEs) at multiple scales, and use Robin to identify shortcomings of current evaluation approaches across scales. Next, to overcome the identified limitations, we introduce CHIRP - a new long form response benchmark we developed for more robust and complete VLM evaluation. We provide open access to the Robin training code, model suite, and CHIRP benchmark to promote reproducibility and advance VLM research.
@@ 1 Introduction
To demonstrate the limitations of existing benchmarks and the efficacy of our proposed method, we introduce Robin, a suite of VLMs trained at various scales, inspired by the Pythia language model suite Biderman et al. (2023). By systematically varying the Vision Encoder (VE) and the Large Language Model (LLM) sizes, we will show that while benchmark scores remain largely unaffected, human evaluations reveal significant differences in the models’ outputs quality.
Our findings underscore the need for more robust and human-centric VLM evaluation methodologies. CHIRP paves the way for developing more reliable and informative VLM benchmarks, ultimately leading to the creation of more effective and impactful VLMs.
Our Contributions:
  • We investigate the drawbacks of relying on automatic metrics and show the benefits of AI-based and human-based evaluations of VLMs.
  • We present CHIRP, an open-ended question-and-answer benchmark.
  • We train and release an open-source collection of VLMs named Robin. Robin is a scaling suite based on LLMs and VEs of different sizes. This allows to study the effects of scaling both language and vision components on downstream performance of VLMs.
  • We compare the performance of the trained VLMs using a wide range of evaluation approaches: automated metrics, AI-based evaluations, and human evaluations.
@@ 2 Related Work
Kaplan et al. Kaplan et al. (2020) pioneered the application of scaling laws to language models, demonstrating a power-law relationships between loss and model size, dataset size, and compute time. This has led to practical applications, such as the Pythia suite Biderman et al. (2023), which comprises of identically trained language models with varying parameter sizes, empirically verifying these scaling laws.
Cherti et al. Cherti et al. (2023) investigated the scaling laws of the CLIP vision encoders, training and comparing different sizes of the CLIP vision encoders on the same data. These models indeed verified the aforementioned scaling laws and have become a very popular suite of models.
@@
The advent of powerful foundation models like GPT-4V offers a new way to evaluate weaker models, moving beyond traditional, rigid metrics such as exact string matching, as done in Hudson & Manning (2019); Mishra et al. (2019); Singh et al. (2019). Early evidence from benchmarks like MM-Vet Yu et al. (2023) and VQA tasks Agrawal et al. (2016) suggests that evaluating with stronger models offers a promising path towards more comprehensive and insightful evaluation, surpassing the limitations of static, string-based methods Ji et al. (2023); Lee et al. (2024). This shift towards leveraging the semantic understanding of LLMs for evaluation promises to unlock a better understanding of model capabilities.
Model Parameter Count
CLIP ViT B 86 million
CLIP ViT L 307 million
CLIP ViT H 632 million
CLIP ViT g 1 billion
@@
Zheng et al. Zheng et al. (2023) introduce two benchmarks, MT-Bench and Chatbot Arena, to explore the feasibility of employing LLMs as judges. Their findings indicate that advanced LLMs, such as GPT-4, closely align with human preferences, achieving over 80% of agreement Rafailov et al. (2024). Similarly, AlpacaEval Li et al. (2023a) utilizes LLMs to assess instruction-following models.
@@
Wu et al. Wu & Aji (2023) focused on the bias in evaluations conducted by both human and LLM annotators, particularly noting a preference for flawed content if it avoids brevity or grammatical errors, and introduced the Multi-Elo Rating System (MERS) for more nuanced assessments. A study by Koo et al. Koo et al. (2023) pointed out significant biases of LLMs evaluators, with an average Rank-Biased Overlap (RBO) score of 49.6%, suggesting a misalignment between machine and human preferences.
3 Robin VLM Suite: Training Methodology
We review the methodology used to train our scaling suite, and the different experiments conducted with the trained models.
3.1 Model Architectures
Our models are based on the LLaVA architecture Liu et al. (2023a; b) and consist of three components: a pretrained vision encoder, a MultiLayer Perceptron (MLP) projection that converts image features to text space, and a language model that uses self-attention to process both visual and textual tokens. Each component can be individually tuned during training. The exact training process, including steps, data composition, and hyperparameters, is detailed in Appendix A.
3.2 Experimental Design
To design a scaling suite for VLMs, we vary the language encoder and vision encoder. Our setup is based on the Pythia suite Biderman et al. (2023), which maintains consistent training data and order, leaving model size as the only variable. Similarly, the CLIP Radford et al. (2021) vision encoders released by LAION follow this pattern. We train VLMs using 5 Pythia sizes (410M, 1.4B, 2.8B, 6.9B, and 12B parameters) paired with 4 CLIP models (Base, Large, Huge, and gigantic). The sizes of these CLIP models are detailed in Table 2, resulting in 20 Robin models. The scaling laws for the Robin suite over VE size and LLM size is shown in Figure 1.
We run experiments across all Robin models or across two main ablations:
  1. 1.
  2. 2.
4 Benchmark Results
Figure 2: Heatmap showing the scaled average score of the different models of the scaling suite (higher is better).
We ran our suite of models on the following benchmarks: ScienceQA Lu et al. (2022), GQA Hudson & Manning (2019), VQAv2 Goyal et al. (2017), TextVQA Singh et al. (2019), MM-Vet Yu et al. (2023), and LLaVA-Bench Liu et al. (2023b). The complete results of the models on these benchmarks are detailed in Appendix A.4, which includes a complete score table (Table 5) and heatmaps for all benchmarks (Figure 10). Figure 2 shows the scaled average scores. Due to varying score distributions across benchmarks, we use a scaled average. For example, VQAv2 scores range from 40 to 60, while MM-Vet scores range from 6 to 18. The scaled score is calculated as follows: let $S$ be the matrix of scores, with each row $S_{i,:}$ representing the scores model $i$ obtained on all $N$ benchmarks, and $S_{:,j}$ representing the scores of all models on benchmark $j$. Let $S^{*}$ be the scaled scores vector.
The scaled average is plotted in Figure 2. As it is shown, there is no clear relationship between VE size and model performance. However, a slight trend between LLM size and performance is observed. Despite this, empirical testing revealed significant differences between the models that these benchmarks did not capture.
5 Investigating Existing Benchmarks
Empirical testing suggested that existing benchmarks might not capture all observed model capabilities. We aimed to determine whether the standard evaluation methods were inaccurate or if the benchmarks themselves were flawed.
In 100 questions sampled from GQA, we found that 9 questions had incorrect ground truth answers. If we want to estimate the error of this value, the actual percentage of incorrect prompts $\hat{p}$ is $\hat{p}\in p\pm z*\sqrt{\frac{p*(1-p)}{n}}$. For a 95% confidence interval: $z=1.96$, and with our sample size $n$ of 100, we measured $p=0.09$, we are 95% certain: $3.4\%\leq\hat{p}\leq 14.6\%$. This equates to 770,769 to 3,309,773 questions of the 22,669,678 GQA questions being incorrect. Although this is a large spread, this result rmeains quite significant, as most improvements on State of The Art (SoTA) models are very small, regularly under 3% Li et al. (2023b). These findings lead to the conclusion that if 2 models score within 3% of each other on GQA, they could very well be equal in actual performance on it. Representative examples of the aforementioned questions are shown in Appendix B.4.1.
Ultimately, after examining benchmarks and responses, we propose the following hypotheses for why our models did not exhibit expected scaling trends:
  • benchmarks were graded inaccurately
  • vague questions with multiple possible answers and incorrect ground truth answers
  • questions themselves don’t demand a detailed examination of images
In the following subsections, we test each of the above hypothesis to see if addressing these issues reveal trends in model scale we hadn’t observed previously.
5.1 Long vs Short Responses (LvSR)
The GQA benchmark provides an evaluation script that grades responses using string matching on single phrase responses. On the sample of 100 GQA questions, we prompted and manually graded our models for LvSR to see if new trends across the LLM size ablation appear with longer responses, the results of which are shown in Figure 3. For sufficiently large models, we did not notice a significant improvement in overall model accuracy. However, models often got different questions correct when responding with LvSR. To show this, we calculated a superscore, in which responses were marked correct if either the long or short response was correct (See Figure 3). The improved results of the superscore indicate that while long and short responses achieve a similar overall accuracy, they tend to be accurate for a different set of questions. This suggests that evaluating long and short responses demonstrate different model skills. In Appendix B.4.3, we’ve included examples of differing responses when prompting for long and short answers.
5.2 Inaccurate Grading and LLM evaluations
Most existing automatic metrics are incapable of evaluating longer responses, and often fail in scenarios where the models being tested do not output the answer in the expected format Hudson & Manning (2019); Singh et al. (2019). For example, models may respond with a synonym for the ground truth, which can cause issues with exact string matching based evaluation. These issues can be especially prevalent with non instruction tuned models, or small scale models.
To address responses that automated evaluations cannot recognize, we utilize a the GPT-4 LLM OpenAI (2023) to evaluate whether a given response matches the correct answer or not. We ran this LLM evaluations on both the long and short responses, using the prompting detailed in Appendix B.3.
5.3 Multiple Possibilities and VLM evaluation
Our empirical analysis revealed that ground truth answers are not always representative of all possible correct answers. In GQA and TextVQA, this issue arises from ambiguous questions that can have multiple valid answers, as shown in Appendix B.4.1 and B.4.2. In questions where ground truth answers don’t encompass all valid answers, LLMs don’t have sufficient information to accurately responsed.
A comparison of the accuracy of LLaVA-34B and GPT-4V over LLM scale can be seen in Figure 4. GPT-4V evaluations of GQA differed from human evaluations more than LLaVA-34B due to GPT-4V applying stricter grading criteria. Appendix B.4.3 presents a few such examples. Although LLaVA-34B had higher accuracy, we hypothesize that further work could align GPT-4V’s grading schema closer to the LLaVA-34B grading by prompting for a looser grading, likely leading to improved results for GQA evaluation.
5.4 Comparison of automated, LLM, and VLM evaluations
We graph scaling across LLM size, and VE size using all AI evaluation methods in Figure 4. AI evaluations seem to yield different and more accurate results across the largest LLM and VE sizes. However, despite the improved accuracy in evaluation techniques, our models still did not exhibit expected scaling relationships, the complete results being shown in Appendix B.1. We hypothesize that long form questions may be more conducive to extracting fine grain estimation of model knowledge than questions from existing benchmarks.
Figure 4: Evaluating GQA and textVQA using automated string matching programs, LLMs, VLMs, and humans evaluators. Solid lines are evaluations of responses where models were prompted for short responses. Dashed lines have no such prompting for brevity. Left. Robin GQA and textVQA scores across LLM size and VE size ablations, calculated using the appropriate evaluation method for each evaluator. Right. Accuracy of those evaluation methods on the GQA and textVQA sample over the LLM size ablation. The accuracy was determined by comparing evaluations to the human grading.
6 CHIRP benchmark
@@
To address the drawbacks of existing benchmarks outlined in Section 5, we introduce CHIRP, a new evaluation benchmark, which grades long form responses. CHIRP comprises of 104 open ended questions, evaluated by either humans or VLMs. These free form questions do not correspond to a single "correct" answer. Instead, they require models to generate flexible, creative and complex responses. Consequently, we evaluate models using a preference based rating in which two model’s responses are compared side by side. Instructions on downloading the CHIRP benchmark can be found in Appendix C.
@@ 6.1 Generating the dataset
We wrote questions along with image descriptions, which we then refined with the help of GPT-4 OpenAI (2023). The image descriptions were given to Dalle-E 3 to generate the associated images. We would then iterate and finetune the description by hand in order to get the desired image.
@@
Unlike many datasets that rely on pre-existing images, our approach allows us to generate images specifically tailored for thought-provoking questions and detailed analysis. This also removes the risk of the model having seen the image in training. Moreover, we eliminate the risk of evaluating models on contaminated images as all of them were validated by hand.
@@ 6.2 Human based evaluations
We utilized CloudResearch for large scale human evaluation of our model’s responses. To this end, we presented users with the responses of two models and asked them to indicate their preferred response on a set of criteria. There are 5 criteria: overall preference, relevance and completeness, understanding and reasoning, hallucinations, and details. These criteria were chosen as empirical evidence showed that these were under-evaluated in other benchmarks and the most important to a user’s perception of the model quality. An example of the user interface as well as a detailed description of each criterion can be found in Appendix C.2.1.
We validated this evaluation method by evaluating our suite of models on all five criteria across the LLM size and VE size ablations. Due to limitations in time and budget, for each question of the dataset, we randomly sample five model matchups out of all the model pairwise combinations. We also ran evaluations across our entire suite of VLMs to judge the overall preference criteria. To this end, we randomly selected 25 matchups from the 190 possible pairs of Robin models. Full details on the human evaluation setup can be found in Appendix section C.2.2.
@@ 6.3 VLM based evaluations
To evaluate our models on CHIRP at scale, we experiment with the use of VLMs: GPT-4V and LLaVA-34B. Rather than asking a human for model preferences, we asked the VLMs to indicate their preferred response for each criterion. For GPT-4V, we utilized two distinct prompts: GPT-4V (S) (simple), which directly solicited model preferences, and GPT-4V (R) (reasoning), which prompted the VLM to reason before making a decision. We extracted the VLMs final choice using GPT-3.5. Detailed explanations of these prompts are provided in Appendix C.2.3.
We evaluated all combinations of matchups from the LLM size and VE size ablations across all criteria. We also ran GPT-4V (R) evaluations on a random sample of 50 matchups from all combinations of Robin models on the overall preference criteria.
Figure 5: Mean Elo calculated over LLM Size (top row) and VE Size (bottom row) using different evaluators (columns) and criteria (series). Graphs are calculated using bootstrapping on 1000 samples. Each sample is drawn with low transparency and the solid lines indicate the mean over samples for the respective category.

### v3

# CHIRP: A Fine-Grained Benchmark for Open-Ended Response Evaluation in Vision-Language Models

@@ Abstract
The proliferation of Vision-Language Models (VLMs) in the past several years calls for rigorous and comprehensive evaluation methods and benchmarks. This work analyzes existing VLM evaluation techniques, including automated metrics, AI-based assessments, and human evaluations across diverse tasks. We test our model on a scaling suite across vision and langauge sizes, identifiying shortcomings of current evaluation approaches. To overcome the identified limitations, we introduce CHIRP - a new long form response benchmark we developed for more robust and complete VLM evaluation. We provide open access to our evaluation code and the CHIRP benchmark to promote reproducibility and advance VLM research.
@@ 1 Introduction
To demonstrate the limitations of existing benchmarks and the efficacy of our proposed method, we use a scaling suite called Robin, a series of VLMs trained at various scales, inspired by the Pythia language model suite Biderman et al. (2023). By systematically varying the Vision Encoder (VE) and the Large Language Model (LLM) sizes, we see that while some benchmark scores remain largely unaffected, human evaluations reveal significant differences in the models’ outputs quality.
Our findings underscore the need for more robust and human-centric VLM evaluation methodologies. CHIRP paves the way for developing more reliable and informative VLM benchmarks, ultimately leading to the creation of more effective and impactful VLMs. Furthermore, we will also investigate in detail the possibility of using different AI agents as proxies to human evaluations.
@@ 2 Related Work
Kaplan et al. (2020) pioneered the application of scaling laws to language models, demonstrating a power-law relationships between loss and model size, dataset size, and compute time. This has led to practical applications, such as the Pythia suite Biderman et al. (2023), which comprises of identically trained language models with varying parameter sizes, empirically verifying these scaling laws.
Cherti et al. (2023) investigated the scaling laws of the CLIP vision encoders, training and comparing different sizes of the CLIP vision encoders on the same data. These models indeed verified the aforementioned scaling laws and have become a very popular suite of models.
@@
The advent of powerful foundation models like GPT-4V offers a new way to evaluate weaker models, moving beyond traditional, rigid metrics such as exact string matching, as done in popular benchmarks like Hudson & Manning (2019); Mishra et al. (2019); Singh et al. (2019). Early evidence from benchmarks like MM-Vet Yu et al. (2023) and VQA tasks Agrawal et al. (2016) suggests that evaluating with stronger models offers a promising path towards more comprehensive and insightful evaluation, surpassing the limitations of static, string-based methods Ji et al. (2023); Lee et al. (2024). This shift towards leveraging the semantic understanding of LLMs for evaluations promises to unlock a better evaluations, leading to a better understanding of model capabilities.
@@
Zheng et al. (2023) introduce two benchmarks, MT-Bench and Chatbot Arena, to explore the feasibility of employing LLMs as judges. Their findings indicate that advanced LLMs, such as GPT-4, closely align with human preferences, achieving over 80% of agreement Rafailov et al. (2024). Similarly, AlpacaEval Li et al. (2023a) utilizes LLMs to assess instruction-following models.
@@
Wu & Aji (2023) focused on the bias in evaluations conducted by both human and LLM annotators, particularly noting a preference for flawed content if it avoids brevity or grammatical errors, and introduced the Multi-Elo Rating System (MERS) for more nuanced assessments. A study by Koo et al. (2023) pointed out significant biases of LLMs evaluators, with an average Rank-Biased Overlap (RBO) score of 49.6%, suggesting a misalignment between machine and human preferences.
While there exists benchmarks like LLaVA-Bench Liu et al. (2023a) which elicit open-ended responses, and evaluate using LLMs, they often lack thought-provoking and diverse questions that challenge a model’s perception and contextual understanding. For example, most of the open-ended questions in LLaVA-Bench are of the form: "describe the image…". We address this shortcoming with our benchmark CHIRP.
3 CHIRP
@@
We introduce CHIRP, a new evaluation benchmark, which grades long form responses. CHIRP comprises of 104 open ended questions, evaluated by either humans or VLMs. These free form questions do not correspond to a single "correct" answer. Instead, they require models to generate flexible, creative and complex responses. Consequently, we evaluate models using a preference based rating in which two model’s responses are compared side by side. Instructions on downloading the CHIRP benchmark can be found in Appendix C.
We will show in Section 5.2 that current benchmarks evaluate models with high variance, making it difficult to accurately rank models with similar performance. CHIRP addresses this issue by using significantly fewer but higher-quality prompts and pairwise evaluations to establish a more reliable ranking.
@@ 3.1 Generating the dataset
We wrote questions along with image descriptions, which were then refined with the help of GPT-4 OpenAI (2023). The image descriptions were given to Dalle-E 3 to generate the associated images. We finetuned the image descriptions and questions by hand to generate the optimal image question pairs.
@@
Unlike many datasets that rely on pre-existing images, our approach allows us to generate images specifically tailored to our thought-provoking questions and detailed analysis. This also removes the risk of the model having seen the image in training. Moreover, we eliminate the risk of evaluating models on contaminated images as all of them were validated by hand. A sample text-image pair of the CHIRP dataset can be seen in Appendix C.2.1.
@@ 3.2 Human based evaluations
We utilize CloudResearch for large scale human evaluation of our model’s responses. To this end, we presente users with the responses of two models to the same question and ask them to indicate their preferred response on a set of criteria. There are 5 criteria: overall preference, relevance and completeness, understanding and reasoning, hallucinations, and details. These criteria were chosen as empirical evidence showed that these were the most important to a user’s perception of the model quality. An example of the user interface as well as a detailed description of each criterion can be found in Appendix C.2.1.
@@ 3.3 VLM based evaluations
To evaluate our models on CHIRP at scale, we experiment with the use of VLMs, namely GPT-4V and LLaVA-34B, as judges. Rather than asking a human for model preferences, we asked the VLMs to indicate their preferred response for each criterion. For GPT-4V, we utilized two distinct prompts: GPT-4V (S) (simple), which directly solicited model preferences, and GPT-4V (R) (reasoning), which prompted the VLM to reason before making a decision. We extracted the VLMs final choice using GPT-3.5. Detailed explanations of these prompts are provided in Appendix C.2.3.
@@ 3.4 Elo ratings
4 Experimental Setup
To evaluate CHIRP’s effectiveness, we test it on a suite of VLMs with controlled single-parameter variations. This setup allows us to assess whether CHIRP can capture subtle differences in model performance that existing benchmarks fail to detect. We outline the details of the models evaluated on and the benchmarks compared against below.
Evaluation models
We test CHIRP on a suite of models based on the LLaVA architecture, with controlled variations in both the LLM size and the Vision Encoder (VE) size. Each model is built by combining an LLM from the Pythia suite with a VE from the CLIP suite. These models, referred to as Robin, include three main experimental groups:
  1. 1.
    All Robin models - 20 models with 5 Pythia sizes (410M, 1.4B, 2.8B, 6.9B, and 12B parameters) paired with 4 CLIP models (Base, Large, Huge, and gigantic)
  2. 2.
  3. 3.
By running experiments across these Robin models, we aim to demonstrate CHIRP’s ability to capture subtle differences in model performances that are often missed by traditional benchmarks but have noticeable differences in response quality. We illustrate this complete pipeline in Figure 1.
More on the architecture and training procedure for these models can be found in Appendix A.
Benchmark Comparisons
We ran our suite of models on the following benchmarks: ScienceQA Lu et al. (2022), GQA Hudson & Manning (2019), VQAv2 Goyal et al. (2017), TextVQA Singh et al. (2019), MM-Vet Yu et al. (2023), and LLaVA-Bench Liu et al. (2023b).
We calculate a scaled average to show the general results from all benchmarks. The scaled score is calculated as follows: let $S$ be the matrix of scores, with each row $S_{i,:}$ representing the scores model $i$ obtained on all $N$ benchmarks, and $S_{:,j}$ representing the scores of all models on benchmark $j$. Let $S^{*}$ be the scaled scores vector.
Running Robin on CHIRP
For human based evaluations, we validate this method by evaluating our suite of models on all five CHIRP criteria across the LLM size and VE size ablations. Due to limitations in time and budget, for each question of the dataset, we randomly sample five model matchups out of all the possible model pairwise combinations. We also ran evaluations across our entire suite of VLMs to judge the overall preference criteria. To this end, we randomly selected 25 matchups from the 190 possible pairs of Robin models. Full details on the human evaluation setup can be found in Appendix section C.2.2.
For AI based evaluations: we evaluate all combinations of matchups from the LLM size and VE size ablations across all criteria. We also run GPT-4V (R) evaluations on a random sample of 50 matchups from all combinations of Robin models on the overall preference criteria.
Figure 1: Experimental Setup Pipeline. Detailed pipeline used to evaluate CHIRP: from the left we have the models trained with the 3 different ablations, then in the middle we evaluate said models on different benchmarks, including CHIRP, and finally we compare the model rankings obtained on the different benchmarks.
5 Evaluation Results
5.1 Comparison with existing benchmarks
In all benchmarks, there is no discernible relationships between VE size and model performance. The scaled average of the results are plotted in Figure 2. As it is shown, there is no clear relationship between VE size and model performance. However, a slight trend between LLM size and performance is observed. The complete results of the models on these benchmarks are detailed in Appendix A.4, which includes a complete score table and heatmaps for all benchmarks. In comparison, a preliminary analysis of the results on CHIRP clearly show the effect of language model scaling.
Figure 2: All Robin model performance on existing benchmarks vs CHIRP. Left. Heatmap showing the scaled average score detailed in Section 4. Center. Elo calculated from GPT-4V (R) on CHIRP. Right. Elo calculated from human survey on CHIRP.
5.2 Investigating differences between CHIRP and other benchmarks
Empirically testing our models we saw that model performance aligned more closely to the trends indicated using CHIRPs evaluations. Here we aim to determine the source of why existing benchmarks might not capture all observed model capabilities, and in doing so highlight the gap in benchmarking that CHIRP hopes to fill.
We hypothesized that existing benchmarks may not capture subtle changes between models for one of the following reasons:
  1. 1.
  2. 2.
    benchmarks aren’t graded with sufficient accuracy
  3. 3.
    questions don’t demand a detailed examination of images or their ground truths may not be accurate
In the following subsections, we test each of the above hypothesis to see if addressing these issues yields results more closely aligned with those of CHIRP and if it could help improve the targeted benchmarks. To do so, we sample 100 questions from both the GQA and textVQA benchmarks.
5.2.1 Long vs Short Responses (LvSR)
The GQA benchmark provides an evaluation script that grades responses using string matching on single phrase responses. On the sample of 100 GQA questions, we prompted and manually graded our models for LvSR to see if new trends across the LLM size ablation appear with longer responses, the results of which are shown in Figure 3. For sufficiently large models, we did not notice a significant improvement in overall model accuracy. However, models often got different questions correct when responding with LvSR. To show this, we calculated a superscore, in which responses were marked correct if either the long or short response was correct (See Figure 3). The improved results of the superscore indicate that while long and short responses achieve a similar overall accuracy, they tend to be accurate for a different set of questions. This suggests that evaluating long and short responses demonstrate different model skills. In Appendix B.4.3, we’ve included examples of differing responses when prompting for long and short answers. We believe that this is one of the reasons CHIRP indicates different evaluation results as compared to existing benchmarks.
5.2.2 Inaccurate Grading
LLM Grading.
Most existing automatic metrics are incapable of evaluating longer responses, and often fail in scenarios where the models being tested do not output the answer in the expected format, as shown in Hudson & Manning (2019); Singh et al. (2019). For example, models may respond with a synonym for the ground truth, which can cause issues with exact string matching based evaluations. These issues can be especially prevalent with non instruction tuned models, or small scale models.
To address responses that automated evaluations cannot recognize, we utilize GPT-4 OpenAI (2023) to evaluate whether a given response matches the correct answer or not. We run this LLM evaluation on both the long and short responses, using the prompting detailed in Appendix B.3.
VLM Grading.
Our empirical analysis revealed that ground truth answers are not always representative of all possible correct answers. In GQA and TextVQA, this issue arises from ambiguous questions that can have multiple valid answers, as shown in Appendix B.4.1 and B.4.2. In questions where ground truth answers don’t encompass all valid answers, LLMs don’t have sufficient information to accurately respond.
Human Grading.
We manually grade the responses generated from models in the LLM Size ablation. We use these grades to calculate the accuracy of the automated, LLM and VLM evaluations. The accuracy is determined by the percentage of responses where the evaluation agrees with the human grading. The accuracy is plotted in the center plots of Figure 4.
Comparison of Evaluation Methods.
We graph scaling across LLM size, and VE size using the automated and AI evaluation methods in Figure 4.
We observe that long-form response accuracy exhibits greater sensitivity to model parameter count. Furthermore, prompting short-form benchmarks for extended responses does not enhance grading accuracy. These findings highlight the necessity of explicitly designing benchmarks for long-form prompting, a challenge we address with CHIRP.
Additionally, our findings indicate that neither AI-based nor traditional automated evaluations show a consistently stronger agreement with human assessments, as evaluation accuracy remains inconsistent across short and long responses, with no method proving definitively superior. This suggests that CHIRP’s improvements stem primarily from its design favoring open-ended questions over those with ground-truths.
Figure 4: Left. Evaluating GQA and textVQA across the LLM size and VE size ablations using automated string matching programs, LLMs, VLMs, and humans evaluators. Solid lines are evaluations of responses where models were prompted for short responses. Dashed lines have no such instruction. Right. Accuracy of these evaluation methods on the GQA and textVQA sample over the LLM size ablation. The accuracy was determined by comparing evaluations to the human grading.
5.2.3 Poor questions and ground truths
In 100 questions sampled from GQA, we found that 9 questions had incorrect ground truth answers. If we want to estimate the error of this value, the actual percentage of incorrect prompts $\hat{p}$ is $\hat{p}\in p\pm z*\sqrt{\frac{p*(1-p)}{n}}$. For a 95% confidence interval: $z=1.96$, and with our sample size $n$ of 100, we measured $p=0.09$, we are 95% certain: $3.4\%\leq\hat{p}\leq 14.6\%$. This equates to 770,769 to 3,309,773 questions of the 22,669,678 GQA questions being incorrect. Although this is a large spread, this result remains quite significant, as most improvements on State of The Art (SoTA) models are very small, regularly under 3% Li et al. (2023b). These findings lead to the conclusion that if 2 models score within 3% of each other on GQA, they could very well be equal in actual performance on it. Representative examples of the aforementioned questions are shown in Appendix B.4.1.
This leads us to our main conclusion for this section. Errors in ground truths lead to inaccurate benchmark scores. Additionally, the fact that most modern SoTA models fall within the margin of error of these benchmarks highlights the need for a precise, pairwise, benchmark like CHIRP.
6 Using CHIRP in practice
We have not yet discussed the practicality of using our benchmark. In general, human evaluations are expensive, and thus are not the standard method for evaluating model performance. Here we investigate the extent to which AI evaluations can be used as a proxy for human evaluations in CHIRP.
Figure 5: Mean Elo calculated over LLM Size (top row) and VE Size (bottom row) using different evaluators (columns) and criteria (series). Graphs are calculated using bootstrapping on 1000 samples.
6.1 Qualitative Analysis - Trends